# Mixed Hodge Modules and Canonical Perverse Extensions for Multi-Node Conifold Degenerations

### 07 Apr 2026 v1

@@ Abstract.
We study one-parameter conifold degenerations whose central fiber has finitely many ordinary double points and construct a mixed-Hodge-module refinement of the canonical corrected perverse object associated with the degeneration. We build a rank-one point-supported mixed-Hodge-module block at each node, identify the global singular quotient as
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
and assemble these local blocks via Saito’s divisor-case gluing formalism into a global object
$\mathcal{P}^{H}\in MHM(X_{0}).$
We prove that $\mathcal{P}^{H}$ realizes the corrected perverse object, fits into an exact sequence
@@
and that the same quotient realizes the finite local vanishing sector in the nearby-cycle formalism. We further relate the mixed-Hodge-module extension, its realized perverse extension, and the induced extension on hypercohomology carrying the limiting mixed Hodge structure. This gives a theorem-level Hodge-theoretic refinement of the corrected perverse extension in the finite multi-node ordinary double point setting.
@@ 1. Introduction
$\pi:\mathcal{X}\to\Delta$
be a one-parameter degeneration of complex algebraic varieties whose central fiber $X_{0}$ has finitely many ordinary double points
@@
Associated with this degeneration is the canonical corrected perverse sheaf $\mathcal{P}$, obtained from the nearby- and vanishing-cycle formalism. In the single-node case, this object was constructed and characterized in [11] and shown to be the unique minimal Verdier self-dual perverse extension of the shifted constant sheaf across the node in [10]. Additionally, in [10], the construction was placed in the context of limiting mixed Hodge theory and Saito’s nearby-cycle formalism for mixed Hodge modules, thereby isolating the main Hodge-theoretic problem left open by the perverse-sheaf picture: the construction of a fully internal mixed-Hodge-module refinement of the corrected perverse object.
The purpose of the present paper is to solve that problem in the case of finitely many ordinary double points. More precisely, we construct an object
@@
whose realization under
$\mathrm{rat}:MHM(X_{0})\to\mathrm{Perv}(X_{0};\mathbb{Q})$
recovers the canonical corrected perverse object $\mathcal{P}$, and whose singular quotient is the finite direct sum of the rank-one point-supported mixed Hodge modules attached to the nodes,
@@
We then show that the same mixed-Hodge-module construction governs the vanishing part of the limiting mixed Hodge structure through nearby-cycle hypercohomology. In this way, the present paper supplies the fully internal Hodge-theoretic refinement that was only identified as a gluing problem in [10].
The central point is that the corrected perverse object is not merely a constructible or perverse avatar of degeneration data, but admits a refinement internal to Saito’s category of mixed Hodge modules. This refinement is the missing link between the canonical perverse extension on the singular fiber and the Hodge-theoretic structure carried by nearby cycles. For a finite set of ordinary double points, the singular contribution is point-supported and rank one at each node, so the global problem becomes one of gluing the local ODP mixed-Hodge-module blocks into a single object on $X_{0}$ compatible with the variation morphism, the finite node-supported quotient, and the limiting mixed Hodge structure.
A second theme of the paper is that the finite multi-node setting carries nontrivial global extension-theoretic structure. The singular quotient is not just a bookkeeping device: it is a finite sum of localized rank-one sectors, one at each node, and the corresponding global corrected extension is organized by nodewise extension data. On the perverse side, this yields a distinguished finite-node extension picture which serves as the first precise mathematical shadow of the later quiver-, schober-, and wall-crossing directions. On the Hodge-theoretic side, the same quotient controls the finite vanishing sector in the limiting mixed Hodge structure. Thus the corrected extension is now visible simultaneously in $MHM(X_{0})$, in $\mathrm{Perv}(X_{0};\mathbb{Q})$, and on hypercohomology in the category of mixed Hodge structures.
@@ 1.1. Relation to earlier work
The starting point for the present paper is the construction carried out in [11]. There the corrected perverse object
$\mathcal{P}$
was defined in the ordinary double point case and shown to satisfy the structural properties expected of a canonical extension: it is perverse, restricts to the shifted constant sheaf on the smooth locus, is Verdier self-dual, and fits into an exact sequence
$0\to IC_{X_{0}}\to\mathcal{P}\to i_{*}\mathbb{Q}_{\{p\}}\to 0$
in the single-node case. Moreover, [10] established three points relevant here: (1) in §6 it proved that $\mathcal{P}$ is the unique minimal Verdier self-dual perverse extension of $\mathbb{Q}_{U}[3]$ across the node; (2) it framed the construction in terms of nearby and vanishing cycles, Picard–Lefschetz monodromy, and spherical-twist/schober-type categorical motivation; and (3) it studied the Hodge-theoretic content of the same perverse construction.
The main result in [10] was not the existence of a full mixed-Hodge-module lift, but rather the identification of a common nearby-cycle origin for the corrected perverse object and the corresponding degeneration data on the Hodge-theoretic side. In particular, we showed that the rank-one singular contribution in the perverse extension and the rank-one vanishing contribution in the limiting mixed Hodge structure arise from the same nearby-cycle and vanishing-cycle formalism, while isolating the explicit mixed-Hodge-module gluing problem as the next theorem to be proved.
The present paper fills precisely that gap. It is therefore neither a repetition of the perverse-sheaf arguments of [11] nor a mere reformulation of the Hodge-theoretic framework of [10]. Its role is to solve the missing construction problem: to build the mixed-Hodge-module refinement $\mathcal{P}^{H}$, prove that
$\mathrm{rat}(\mathcal{P}^{H})\cong\mathcal{P},$
establish the exact sequence in $MHM(X_{0})$, and identify the point-supported quotient with the finite vanishing part of the degeneration on the Hodge-theoretic side. In addition, the present paper makes explicit the finite-node extension-theoretic structure of the corrected object and its organization across the three levels $MHM$, perverse sheaves, and mixed Hodge structures.
@@ 1.3. Physical and categorical motivation
Although the present paper is purely mathematical in its statements and proofs, the motivating geometry comes from the conifold transition picture in Calabi–Yau threefolds. In the ordinary double point case, the degeneration is governed by a collapsing three-sphere and the resulting rank-one Picard–Lefschetz monodromy on middle homology. In the physical interpretation of Strominger [19], this collapse corresponds to an additional light BPS state localized at the singularity. In the multi-node case, one expects a finite collection of such localized sectors, one for each node, together with global coupling data reflecting how the local contributions assemble into a single degeneration.
From the categorical side, the rank-one monodromy phenomena of an ODP are mirrored by the rank-one action of spherical twists and, more broadly, by perverse schober structures [17, 5]. The canonical corrected perverse object constructed in [11, 10] may therefore be regarded as the decategorified shadow of a local categorical monodromy phenomenon. The present paper adds the Hodge-theoretic layer needed to refine that picture: it identifies the mixed-Hodge-module object whose realization is the corrected perverse extension, whose quotient isolates the corresponding local vanishing sector on the Hodge-theoretic side, and whose global extension structure already exhibits the formal shape of a finite family of localized sectors coupled to a bulk geometric sector.
@@ 1.4. Main results
Taken together, these results provide the fully internal mixed-Hodge-module refinement of the canonical perverse extension in the finite multi-node ordinary double point case. They also make visible the nodewise extension-theoretic structure of the finite degeneration and furnish the precise Hodge-theoretic foundation needed for subsequent applications to global extension data, quiver-type structures, and the mathematical formulation of localized sectors in conifold-type physical models.
@@ 1.5. Proof strategy
The proof proceeds in four stages. We begin with the local ordinary double point model and construct the mixed-Hodge-module building block that refines the rank-one corrected perverse extension at a single node. This local step isolates the essential nearby-cycle, vanishing-cycle, and divisor-gluing data in Saito’s formalism and identifies the point-supported mixed Hodge module that refines the local vanishing contribution. The local construction is then used as the atomic input for the finite multi-node case.
The remaining stages pass from local to global. First, we prove that for finitely many ordinary double points the singular quotient is supported on the node set and decomposes into one rank-one point-supported mixed Hodge module per node. We then analyze the corresponding extension spaces and show that the finite-node corrected extension admits a nodewise organization. Second, we construct the global object $\mathcal{P}^{H}\in MHM(X_{0})$ by gluing the smooth-locus Hodge module to the finite collection of local node contributions and prove that its realization is the canonical corrected perverse object $\mathcal{P}$. Finally, we analyze the hypercohomology of $\mathcal{P}^{H}$ and show that the same nearby-cycle mixed-Hodge-module formalism governs the vanishing part of the limiting mixed Hodge structure. In this way, the local ODP model, the global extension on $X_{0}$, and the Hodge-theoretic degeneration data are placed in a single framework.
@@ 3.3. Saito’s divisor gluing theorem
The formal engine of the present paper is Saito’s gluing theorem for a principal divisor [13, Prop. 0.3]. Let $g:X\to\mathbb{C}$ be a regular function, and let
@@ 4.5. Normalization by monodromy
By Propositions 4.6 and 4.8, both $u_{\mathrm{loc}}$ and $v_{\mathrm{loc}}$ are nonzero and unique up to scalar on the one-dimensional ODP sector. Hence the composite $v_{\mathrm{loc}}u_{\mathrm{loc}}$ differs from the induced action of NN on that sector by a nonzero scalar. Rescaling one of the two maps removes that scalar. ∎
@@ 4.6. Local gluing datum
This follows immediately from Definition 4.4 and Propositions 4.8, 4.9. ∎
@@ 4.7. The local mixed-Hodge-module correction block
Apply Saito’s local gluing formalism to the datum of Theorem 4.10. The hypotheses are satisfied by Theorem 4.10, with bulk term $IC^{H}_{\mathrm{loc}}$ and singular term $W^{H}_{\mathrm{loc}}$, so the gluing theorem produces the required extension object. ∎
@@ 5.6. Realization and compatibility with the perverse extension space
Remark 5.4.
The decomposition above gives a mathematically precise version of the later domain-wall picture: the point-supported node sectors are localized degrees of freedom, while the extension classes in the nodewise summands measure their couplings to the bulk intersection-complex sector.

### 09 Apr 2026 v2

@@ Abstract.
We study one-parameter conifold degenerations whose central fiber has finitely many ordinary double points and construct a mixed-Hodge-module refinement of the canonical corrected perverse object associated with the degeneration. At each node we obtain a rank-one point-supported mixed Hodge module, and globally we identify the singular quotient as
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Using Saito’s divisor-case gluing formalism, we construct an object
$\mathcal{P}^{H}\in MHM(X_{0})$
fitting into an exact sequence
@@
and show that its realization is the corrected perverse object. We then relate the same quotient to the finite local vanishing sector in the nearby-cycle formalism and compare the mixed-Hodge-module extension, its realized perverse extension, and the induced extension on hypercohomology carrying the limiting mixed Hodge structure.
@@ 1. Introduction
$\pi:X\to\Delta$
be a one-parameter degeneration whose central fiber $X_{0}$ has finitely many ordinary double points
@@
Associated with this degeneration is the canonical corrected perverse object $\mathcal{P}$, obtained from the nearby-/vanishing-cycle formalism. In the single-node case, this object was constructed and analyzed in [10], and subsequent work identified the corresponding Hodge-theoretic gluing problem inside Saito’s nearby-cycle formalism [9].
The purpose of the present paper is to solve that gluing problem in the finite multi-node case. We construct an object
@@
whose realization is the corrected perverse object and whose singular quotient is the finite direct sum
@@
We then show that the same quotient governs the finite local vanishing sector on the Hodge-theoretic side through nearby-cycle hypercohomology.
Thus the corrected extension is visible simultaneously at three levels: in $MHM(X_{0})$, after realization in $Perv(X_{0};\mathbb{Q})$, and on hypercohomology in the category of mixed Hodge structures. The finite multi-node case is the first setting in which the local ordinary-double-point blocks, the global extension problem, and the limiting mixed Hodge structure can all be organized in a single mixed-Hodge-module framework.
@@ 1.1. Relation to earlier work
The starting point for the present paper is the single-node construction of the corrected perverse object. In that setting, the corrected object was shown to be perverse, to restrict to the shifted constant sheaf on the smooth locus, and to fit into the expected point-supported extension sequence. Subsequent work recast that construction in terms of nearby and vanishing cycles and compared the perverse correction with the corresponding degeneration data on the Hodge-theoretic side. What was isolated there, but not constructed, was the internal mixed-Hodge-module object whose realization is the corrected perverse extension.
The role of the present paper is to construct that object in the finite multi-node ordinary double point case. Concretely, we build the mixed-Hodge-module refinement $\mathcal{P}^{H}$, identify its singular quotient, prove the corresponding exact sequence in $MHM(X_{0})$, show that realization recovers the corrected perverse object, and relate the same quotient to the finite local vanishing sector on hypercohomology. Accordingly, the present paper is neither a repetition of the earlier perverse-sheaf construction nor a restatement of the nearby-cycle bridge. Its contribution is the mixed-Hodge-module construction itself and the resulting comparison framework among mixed Hodge modules, perverse sheaves, and mixed Hodge structures.
@@ 1.3. Physical and categorical motivation
Although the present paper is purely mathematical in its statements and proofs, its motivating geometry comes from conifold degenerations of Calabi–Yau threefolds. In the ordinary double point case, the local degeneration is governed by a collapsing three-sphere and the corresponding rank-one Picard–Lefschetz monodromy on middle homology. In the multi-node case, one obtains a finite collection of such local vanishing sectors, one at each node.
From the categorical side, rank-one monodromy phenomena of an ordinary double point are closely related to spherical twists and, more broadly, to perverse-schober-type structures. The corrected perverse object constructed in the single-node case may therefore be viewed as a decategorified shadow of a local categorical monodromy phenomenon.
The role of the present paper is not to construct those higher-categorical structures, but to supply the mixed-Hodge-module object that refines the corrected perverse extension and identifies the same finite node-supported quotient on the Hodge-theoretic side. This provides the precise framework needed for later questions about nodewise extension data, limiting mixed Hodge structures, and possible categorical refinements.
@@ 1.4. Main results
Taken together, these results construct the mixed-Hodge-module refinement of the corrected perverse extension in the finite multi-node ordinary double point case, identify its singular quotient, and relate the same quotient to the finite local vanishing sector on hypercohomology. They also isolate the nodewise extension-theoretic structure that underlies the later directions discussed at the end of the paper.
@@ 1.5. Proof strategy
The proof has four stages.
First, we work in the local ordinary double point model and construct the rank-one point-supported mixed-Hodge-module block attached to a single node. This identifies the local singular term and the corresponding gluing morphisms inside Saito’s nearby-cycle formalism.
Second, we pass to the finite multi-node setting and show that the singular quotient is the finite direct sum
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
We then analyze the corresponding extension spaces and isolate the nodewise structure of the finite corrected extension.
Third, we assemble the local gluing data into a global divisor-case gluing datum on $X_{0}$. Saito’s gluing theorem then produces the global mixed Hodge module
$\mathcal{P}^{H}\in MHM(X_{0}),$
and realization identifies its underlying perverse sheaf with the canonical corrected perverse object $\mathcal{P}$.
Finally, we apply hypercohomology to relate the global mixed-Hodge-module extension to the limiting mixed Hodge structure. This identifies the same point-supported quotient with the finite local vanishing sector in the nearby-cycle formalism and yields the comparison among the mixed-Hodge-module extension, its realized perverse extension, and the induced extension on hypercohomology.
@@ 3.3. Saito’s divisor gluing theorem
The formal engine of the present paper is Saito’s gluing theorem for a principal divisor [12, Prop. 0.3]. Let $g:X\to\mathbb{C}$ be a regular function, and let
@@ 4.5. Normalization by monodromy
By Propositions 4.6 and 4.8, both $u_{\mathrm{loc}}$ and $v_{\mathrm{loc}}$ are nonzero and unique up to scalar on the one-dimensional ODP sector. Hence the composite $v_{\mathrm{loc}}u_{\mathrm{loc}}$ differs from the induced action of $N$ on that sector by a nonzero scalar. Rescaling one of the two maps removes that scalar. ∎
@@ 4.6. Local gluing datum
Lemma 4.11 (Filtered structure on the point-supported ODP block).
Let
$W^{H}_{\mathrm{loc}}:=i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1).$
Then:
  1. (1)
    $W^{H}_{\mathrm{loc}}$ is a pure Hodge module supported at $p$.
  2. (2)
    Its underlying filtered $\mathcal{D}$-module is the standard point-supported delta module with the Tate-twisted Hodge filtration.
  3. (3)
    The induced Kashiwara–Malgrange $V$-filtration on $W^{H}_{\mathrm{loc}}$ is the standard point-supported $V$-filtration. In particular, $W^{H}_{\mathrm{loc}}$ is an admissible target for the unipotent nearby-/vanishing-cycle morphisms appearing in Saito’s divisor-case gluing formalism.
  4. (4)
    The morphisms landing in or originating from $W^{H}_{\mathrm{loc}}$ that are induced from the unipotent nearby-/vanishing-cycle formalism are morphisms in the filtered $(\mathcal{D},V)$-module setting.
@@
Since $\mathbb{Q}^{H}_{\{p\}}$ is the constant pure Hodge module on a point, its Tate twist $\mathbb{Q}^{H}_{\{p\}}(-1)$ is again a pure Hodge module on a point. Pushing forward along the closed immersion
$i_{\mathrm{loc}}:\{p\}\hookrightarrow X_{\mathrm{loc}}$
gives
$W^{H}_{\mathrm{loc}}=i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1)\in MHM(X_{\mathrm{loc}}),$
which proves (1).
Its underlying filtered $\mathcal{D}$-module is the standard point-supported delta module with the induced Tate-twisted Hodge filtration, proving (2). For (3), let $t$ be a local defining function for a smooth hypersurface germ through $p$. On the point-supported delta module $\delta_{p}$, one has
$t\cdot\delta_{p}=0$
in the local $\mathcal{D}$-module model. In the local ring $\mathcal{O}_{X,p}$, the delta module $\delta_{p}$ satisfies $t\cdot\delta_{p}=0$, which forces $t\partial_{t}\cdot\delta_{p}=0$ and places $\delta_{p}$ in $V^{1}$ with $t\partial_{t}$-eigenvalue $1$.
Consequently the Kashiwara–Malgrange filtration is the standard point-supported $V$-filtration: the module is concentrated in the single admissible $V$-degree for the unipotent nearby-cycle formalism, equivalently in the degree characterized by the action of $t\partial_{t}$ on the point-supported block; compare the standard local description of the $V$-filtration for point-supported regular holonomic $\mathcal{D}$-modules. This proves (3).
For (4), the canonical and variation morphisms in Saito’s nearby-/vanishing-cycle formalism are morphisms in the filtered $(\mathcal{D},V)$-module setting by construction. Therefore any morphisms induced from them on the local ordinary-double-point sector, in particular the maps used to define $u_{\mathrm{loc}}$ and $v_{\mathrm{loc}}$, remain morphisms in that setting. ∎
Proposition 4.12 (Filtered admissibility of the local ODP gluing datum).
The local datum
is admissible for Saito’s divisor-case gluing formalism. More precisely:
  1. (1)
    $\mathbb{Q}^{H}_{U_{\mathrm{loc}}}[3]$ is a mixed Hodge module on the smooth locus $U_{\mathrm{loc}}$;
  2. (2)
    $W^{H}_{\mathrm{loc}}$ is a point-supported mixed Hodge module with the standard filtered $(\mathcal{D},V)$-module structure of Lemma 4.11;
  3. (3)
    the morphisms $u_{\mathrm{loc}}$ and $v_{\mathrm{loc}}$ are morphisms in the filtered $(\mathcal{D},V)$-module setting and are strict with respect to the Hodge filtration; and
  4. (4)
    the relation
    $v_{\mathrm{loc}}u_{\mathrm{loc}}=N$
    holds on the local ordinary-double-point sector in the filtered setting required by Saito’s gluing theorem.
Proof.
Part (1) is immediate since $U_{\mathrm{loc}}$ is smooth. Part (2) is Lemma 4.11.
For part (3), the morphisms $u_{\mathrm{loc}}$ and $v_{\mathrm{loc}}$ are induced from the unipotent nearby-/vanishing-cycle formalism on the one-dimensional ordinary-double-point sector. The canonical and variation morphisms in that formalism are morphisms in the filtered $(\mathcal{D},V)$-module setting by construction, and morphisms in $MHM$ are strict with respect to the Hodge filtration. It therefore remains to note the $V$-compatibility of the quotient step defining $u_{\mathrm{loc}}$. By construction,
$u_{\mathrm{loc}}=q_{\mathrm{loc}}\circ\mathrm{can}^{H},$
where $q_{\mathrm{loc}}$ extracts the one-dimensional ordinary-double-point vanishing line inside the local vanishing-cycle block. On the local ODP sector this vanishing line is the distinguished one-dimensional summand singled out by the nearby-/vanishing-cycle formalism, hence the quotient map is compatible with the induced $V$-filtration on that sector. Since the ordinary-double-point vanishing line is a direct summand of $\varphi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})$, the projection $q_{\mathrm{loc}}$ onto it is strict with respect to the V-filtration, and therefore the composite $u_{\mathrm{loc}}=q_{\mathrm{loc}}\circ\mathrm{can}^{H}$ is strictly V-compatible. Therefore $u_{\mathrm{loc}}$ is strictly $V$-compatible, and the same argument applies to $v_{\mathrm{loc}}$.
For part (4), the nearby-cycle formalism gives
$\mathrm{Var}^{H}\circ\mathrm{can}^{H}=N$
on the unipotent nearby-cycle object. By the construction of $u_{\mathrm{loc}}$ and $v_{\mathrm{loc}}$ from the local ordinary-double-point sector, together with the normalization of Proposition 4.9, the same relation becomes
$v_{\mathrm{loc}}u_{\mathrm{loc}}=N$
on that sector in the filtered setting. Therefore the local datum satisfies the hypotheses of Saito’s divisor-case gluing theorem. ∎
Remark 4.2 (Scope of the $V$-filtration verification).
The verification in Lemma 4.11 and Proposition 4.12 addresses the specific case of a Tate-twisted point-supported module arising from the one-dimensional ordinary-double-point vanishing sector. In general, Saito’s divisor-case gluing theorem requires filtered compatibility for arbitrary mixed Hodge modules along a divisor, and verifying this may require substantially more information about the underlying filtered $\mathcal{D}$-module structure. In the present setting, however, the point-supported nature of $W^{H}_{\mathrm{loc}}$ reduces the $V$-filtration issue to the standard point-supported case, which is why an explicit local verification is possible here.
@@ 4.7. The local mixed-Hodge-module correction block
By Proposition 4.12, the local datum
satisfies the filtered compatibility hypotheses required by Saito’s divisor-case gluing theorem [12, Prop. 0.3]. Hence that theorem yields an object
$\mathcal{P}^{H}_{\mathrm{loc}}\in MHM(X_{0,\mathrm{loc}})$
whose restriction to the smooth local locus is
$\mathcal{P}^{H}_{\mathrm{loc}}\big|_{U_{\mathrm{loc}}}\cong\mathbb{Q}^{H}_{U_{\mathrm{loc}}}[3]$
and whose singular quotient is
$W^{H}_{\mathrm{loc}}=i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1).$
The kernel of the quotient morphism is the middle extension of the open-part object. Since the middle extension of $\mathbb{Q}^{H}_{U_{\mathrm{loc}}}[3]$ across the isolated local node is $IC^{H}_{\mathrm{loc}}$, we obtain a short exact sequence
$0\to IC^{H}_{\mathrm{loc}}\to\mathcal{P}^{H}_{\mathrm{loc}}\to W^{H}_{\mathrm{loc}}\to 0.$
This is exactly the required local mixed-Hodge-module extension. ∎
@@ 6.2. Assembly of the local gluing morphisms
Proposition 6.4 (Filtered admissibility of the global nodewise gluing datum).
The global datum
is admissible for Saito’s divisor-case gluing formalism. More precisely:
  1. (1)
    $\mathcal{M}^{\prime}_{U}=\mathbb{Q}^{H}_{U}[3]$ is a mixed Hodge module on the smooth locus $U$;
  2. (2)
    is a mixed Hodge module on $X_{0}$ whose filtered structure is obtained componentwise from the local point-supported blocks of Lemma 4.11;
  3. (3)
    the assembled morphisms $u_{\Sigma}$ and $v_{\Sigma}$ are morphisms in the filtered $(\mathcal{D},V)$-module setting and are strict with respect to the Hodge filtration; and
  4. (4)
    the relation
    $v_{\Sigma}u_{\Sigma}=N$
    holds on the finite ordinary-double-point sector in the filtered setting required by Saito’s gluing theorem.
Proof.
Part (1) is immediate. For part (2), $\mathcal{M}^{\prime\prime}_{\Sigma}$ is the finite direct sum of the point-supported Tate-twisted modules treated locally in Lemma 4.11. Since the nodes are pairwise disjoint, the filtered structure on the direct sum is computed componentwise.
For part (3), the global maps $u_{\Sigma}$ and $v_{\Sigma}$ are assembled from the local maps $\{u_{k}\}$ and $\{v_{k}\}$ by Proposition 6.2. Each local pair satisfies the filtered admissibility statement of Proposition 4.12, including compatibility with the induced $V$-filtration on the local ordinary-double-point sector. Finite direct sums preserve strictness and filtered compatibility in this setting, so the assembled maps remain morphisms in the filtered $(\mathcal{D},V)$-module setting and are strict with respect to the Hodge filtration.
For part (4), the identity
$v_{k}u_{k}=N$
holds on each local ordinary-double-point sector by Proposition 4.12. Summing over the finite node set gives
$v_{\Sigma}u_{\Sigma}=N$
on the finite node-supported sector. Therefore the global datum satisfies the hypotheses of Saito’s divisor-case gluing theorem. ∎